Differentiate the following functions with respect to x:
(i) ln(1 + sin² x)
(ii) x ͯ
1
Expert's answer
2020-07-01T19:05:51-0400

i. "ln[1+sin^2x]"

Differentiating with respect to "x:"

"d[ln[1+sin^2x]]\/dx"

"=(1\/(1+sin^2x))*[d(1+sin^2x)\/dx]"

"=" "(1\/(1+sin^2x))*[d[1+(1-cos2x)\/2]\/dx]"

"=(1\/(1+sin^2x))*[d(2+1-cos2x)\/2]\/dx]"

"=(1\/2(1+sin^2x))*[2sin2x]"

"=sin2x\/(1+sin^2x)" (Answer)


ii. "x^x"

Let A="x^x"

Taking "ln" on both sides:

"ln A=xlnx"

Differentiating w.r.t "x" :

"d[ln A]\/dx=d[xlnx]\/dx"

"=(1\/A)*d[A]\/dx=x*(1\/x)+lnx"

or,"dA\/dx=(1+lnx)*A"

"=(1+lnx)*x^x" (Answer)